if the tangent from a pt P on the circle x^2 + y^2 = 1 is perpendicular to tangent frm P to the circle x^2 + y^2 = 3, then locus of P is a circle of radius

a)2 b)3 c)4 d)none of these

Grade:12

2 Answers

bharat bajaj IIT Delhi
10 years ago
THe tangent to circle 1 : x.x1 + y.y1 = 1 from point(x1,y1)
The tangent to circle 2 : x.x1 + y.y1 = 3 from point (x1,y1)
These tangents are perpendicular. Hence,
-x1/y1. -x1/y1 = -1
x1^2 / y1^2 = -1
x1^2 + y1^2 = 0
Hence, NONE OF THESE
THanks
Bharat Bajaj
IIT Delhi
askiitians faculty
vaanya
10 years ago
thanku sir